How long would it take a train traveling 110 mph to get from downtown NY City to downtown Chicago?
Answer by fcabanski(1391)

D=rt (distance = rate times time).


The driving distance is about 790 miles. You'd have to know the rail distance for a more precise answer.


t=D/r ---> t = 790/110 = 7.18 hours or about 7 hours 11 minutes.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)